OpenStudy (anonymous):

A card is drawn from a standard deck of cards. Find the probability that it is a 4 or a red card.

OpenStudy (crashonce):

the standard deck is 52 cards

OpenStudy (crashonce):

consider the red cards, they will be hearts or diamonds, so half of the 52 will be read

OpenStudy (kittiwitti1):

info: -A standard deck of cards has 52 cards -There are four suits that count up from ace to King, so 10 + 3 would make 13 cards per suit. -There are two red suits (hearts and diamonds) and two black suits (clubs and spades). -There is one chance of getting 4 in one suit. conclusion: -red suits: 1/2 chance (2 suits out of 4) -the number 4: once per suit * four suits = 4 chances out of 52 cards.

OpenStudy (kittiwitti1):

._.;

OpenStudy (crashonce):

there are 13 types of cards, so the chance of a 4 is simply 1/13

OpenStudy (kittiwitti1):

Okay maybe I shouldn't have spent two minutes typing that up.

OpenStudy (crashonce):

because the events are independent, we can have 1/2x1/13 = 1/26 am i right?

OpenStudy (kittiwitti1):

...

OpenStudy (anonymous):

Okay, yes, I believe I am following (thank you both very much for the efforts given)

OpenStudy (kittiwitti1):

@CrashOnce is my info incorrect? ._.; (I'm asking seriously.)

OpenStudy (crashonce):

no worries

OpenStudy (kittiwitti1):

You're welcome. ^^;

OpenStudy (crashonce):

no its right

OpenStudy (kittiwitti1):

Okay... I thought you were correcting me there but I wasn't sure xD

OpenStudy (kittiwitti1):

@roseflutterwings do you still need help? :o

OpenStudy (anonymous):

uhm so would the final answer be 1/26 because it is an independent event? Is that all?

OpenStudy (crashonce):

i believe so

OpenStudy (anonymous):

Okay, great :) I can see it. Thank you both!

OpenStudy (crashonce):

no probs

OpenStudy (kittiwitti1):

^-^

OpenStudy (perl):

You can solve this by looking at the 52 card sample space and counting the sample points by hand. Or you can use the formula P( A or B ) = P(A) + P(B) - P(A & B ) Therefore... P( 4 or red card) = P(drawing a 4) + P( drawing red card) - P( drawing four & red) = 4/52 + 26/52 - 2/52